# Solving General-Utility Markov Decision Processes in the Single-Trial Regime with Online Planning

### v1

@@ B.4 Proof of Proposition 2
There exists a one-to-one mapping between histories $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{l}$ in $\mathcal{M}_{f}$, with $10\leq l\leq H-$ 1, and states $\{s,{\boldsymbol{o}}\}\in\mathcal{S}\times\mathcal{O}$ in $\mathcal{M}_{\text{O}}$.
@@
For a given history $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{l}$ in $\mathcal{M}_{f}$, with $10\leq l\leq H-$ 1, consider the mapping defined below that associates $h_{l}$ to a given state $\{s,{\boldsymbol{o}}\}\in\mathcal{S}\times\mathcal{O}$ for $\mathcal{M}_{\text{O}}$ by letting
@@
We aim to show that the mapping above is a bijection between the set of possible histories in $\mathcal{M}_{f}$ and the discrete state space $\mathcal{O}$ in $\mathcal{M}_{\text{O}}$. Clearly, from the mapping above defined, each history $h_{l}$ in $\mathcal{M}_{f}$ is associated with a unique state in $\mathcal{M}_{\text{O}}$. Thus, what remains is to show that any two states $\{s_{1},{\boldsymbol{o}}_{1}\}$ and $\{s_{2},{\boldsymbol{o}}_{2}\}$ for $\mathcal{M}_{\text{O}}$ are equal under mapping (15) if and only if their associated histories $h^{1}$ and $h^{2}$ are equal. We now make two observations. First, for a given state $\{s,{\boldsymbol{o}}\}$ in $\mathcal{M}_{\text{O}}$, component $s$ is directly related, through mapping (15), to the last state in the history $h_{l}$. Second, each history $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})$ will yield through mapping (15) a running occupancy ${\boldsymbol{o}}$ satisfying $\sum_{s,a}o(s,a)=\frac{1-\gamma^{l}}{1-\gamma}$; thus, histories $h_{l}$ with different lengths will yield different ${\boldsymbol{o}}$-vectors. Hence, we only need to show that two running occupancies ${\boldsymbol{o}}_{1}$ and ${\boldsymbol{o}}_{2}$, associated with histories $h^{1}$ and $h^{2}$ (both of length $l$), respectively, are the same if and only if their histories up to timestep $l-1$ are the same:
  • If two histories $h^{1}$ and $h^{2}$ of length $l$ are the same, then it should be clear that their respective running occupancies, as defined through (15), are also the same.
  • If two running occupancies ${\boldsymbol{o}}_{1}$ and ${\boldsymbol{o}}_{2}$ are the same, then their associated histories are also the same. To prove this implication, we focus our attention to a given entry $(s,a)$ of the vectors ${\boldsymbol{o}}_{1}$ and ${\boldsymbol{o}}_{2}$. Running occupancy ${\boldsymbol{o}}_{1}$ is associated with an arbitrary history $h^{1}=(s_{0}^{1},a_{0}^{1},s_{1}^{1},a_{1}^{1},\ldots,s_{l}^{1})$; running occupancy ${\boldsymbol{o}}_{2}$ is associated with an arbitrary history $h^{2}=(s_{0}^{2},a_{0}^{2},s_{1}^{2},a_{1}^{2},\ldots,s_{l}^{2})$. If ${\boldsymbol{o}}_{1}={\boldsymbol{o}}_{2}$ then, for any $s\in\mathcal{S},a\in\mathcal{A}$,
@@
Thus, we conclude that there exists a one-to-one mapping, as defined in (15), between every possible history in $\mathcal{M}_{f}$ up to timestep $H-1$ and states in $\mathcal{M}_{\text{O}}$. ∎
@@ B.5 Proof of Theorem 2
The problem of finding a policy $\pi\in\Pi_{\textnormal{NM}}$ satisfying $\mathcal{R}_{H}(\pi)\leq\epsilon$, for any $\epsilon\in\mathbb{R}_{0}^{+}$, can be reduced to the problem of finding a policy $\pi_{\text{O}}\in\Pi_{\textnormal{S}}$ satisfying $J_{\text{O}}(\pi_{\text{O}})-J_{\text{O}}^{*}\leq\epsilon$. In particular, if $\pi^{*}_{\text{O}}=\operatorname*{arg\,min}_{\pi_{\text{O}}\in\Pi_{\textnormal{S}}}J_{\text{O}}(\pi_{\text{O}})$, then the corresponding non-Markovian policy $\pi$ in $\mathcal{M}_{f}$ satisfies $\mathcal{R}_{H}(\pi)=0$. Finally, it holds that $\mathcal{R}_{H}(\pi)=J_{\text{O}}(\pi_{\text{O}})-J^{*}_{\text{O}}$, where $\pi_{O}$ is the stationary policy for $\mathcal{M}_{\text{O}}$ associated with the non-Markovian policy $\pi$ for $\mathcal{M}_{f}$.
@@
for $F_{1,H}(\pi)$ as defined in (6) and $J_{\text{O}}(\pi_{\text{O}})$ as defined in (9), where $\pi_{\text{O}}$ is the stationary policy for $\mathcal{M}_{\text{O}}$ associated with the non-Markovian policy $\pi$ for $\mathcal{M}_{f}$.
@@
For any $H\in\mathbb{N}$, finite-horizon random trajectories $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots,\mathrm{s}_{H-1},\mathrm{a}_{H-1})$ in $\mathcal{M}_{f}$ are associated with the probability space $(\Omega,\mathcal{F},\mathbb{P}_{\pi})$. We write specific trajectories as $\omega\in\Omega$, with $\omega=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H-1},a_{H-1})$. We highlight that the probability of a given trajectory $\omega\in\Omega$ under policy $\pi\in\Pi_{\text{NM}}$ can be calculated as $\mathbb{P}_{\pi}\left[\omega\right]=p_{0}(s_{0})\cdot\pi(a_{0}|h_{0})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\pi(a_{1}|h_{1})\cdot P^{a_{1}}(s_{1},s_{2})\ldots P^{a_{H-2}}(s_{H-2},s_{H-1})\cdot\pi(a_{H-1}|h_{H-1})$. On the other hand, random trajectories ($(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\{\mathrm{s}_{1},\mathbf{o}_{1}\},\mathrm{a}_{1},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ in $\mathcal{M}_{\text{O}}$ are associated with probability space $(\Omega_{\text{O}},\mathcal{F}_{\text{O}},\mathbb{P}_{\pi_{\text{O}}}^{O})$. We write specific trajectories as $\omega_{\text{O}}\in\Omega_{\text{O}}$, with $\omega_{\text{O}}=(\{s_{0},{\boldsymbol{o}}_{0}\},a_{0},\{s_{1},{\boldsymbol{o}}_{1}\},a_{1},\ldots,\{s_{H},{\boldsymbol{o}}_{H}\})$.
We start by noting that, for any trajectory $\omega_{\text{O}}=(\{s_{0},{\boldsymbol{o}}_{0}\},a_{0},\{s_{1},{\boldsymbol{o}}_{1}\},a_{1},\ldots,\{s_{H},{\boldsymbol{o}}_{H}\})\in\Omega_{\text{O}}$,
$\mathbb{P}_{\pi_{\text{O}}}^{\text{O}}[\omega_{\text{O}}]$ $=p_{0,\text{O}}(\{s_{0},{\boldsymbol{o}}_{0}\})\cdot\pi_{\text{O}}(a_{0}|\{s_{0},{\boldsymbol{o}}_{0}\})\cdot P_{\text{O}}^{a_{0}}(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\{\mathrm{s}_{1},\mathbf{o}_{1}\})\cdot\ldots$
$\qquad\qquad\cdot\pi_{\text{O}}(a_{H-1}|\{s_{H-1},{\boldsymbol{o}}_{H-1}\})\cdot P_{\text{O}}^{a_{H-1}}(\{s_{H-1},{\boldsymbol{o}}_{H-1}\},\{s_{H},{\boldsymbol{o}}_{H}\}).$
=$\overset{(a)}{=}p_{0}(s_{0})\cdot\mathbf{1}({\boldsymbol{o}}_{0}=[0,\ldots,0])\cdot\pi_{\text{O}}(a_{0}|\{s_{0},{\boldsymbol{o}}_{0}\})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\mathbf{1}({\boldsymbol{o}}_{1}=\sigma(s_{0},{\boldsymbol{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\pi_{\text{O}}(a_{H-1}|\{s_{H-1},{\boldsymbol{o}}_{H-1}\})\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\boldsymbol{o}}_{H}=\sigma(s_{H-1},{\boldsymbol{o}}_{H-1},a_{H-1}))$
=$\overset{(b)}{=}p_{0}(s_{0})\cdot\mathbf{1}({\boldsymbol{o}}_{0}=[0,\ldots,0])\cdot\pi(a_{0}|h_{0})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\mathbf{1}({\boldsymbol{o}}_{1}=\sigma(s_{0},{\boldsymbol{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\pi(a_{H-1}|h_{H-1})\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\boldsymbol{o}}_{H}=\sigma(s_{H-1},{\boldsymbol{o}}_{H-1},a_{H-1}))$
@@
=$\overset{(c)}{=}\mathbb{P}_{\pi}\left[\omega\right]\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\boldsymbol{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\boldsymbol{o}}_{1}=\sigma(s_{0},{\boldsymbol{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\mathbf{1}({\boldsymbol{o}}_{H}=\sigma(s_{H-1},{\boldsymbol{o}}_{H-1},a_{H-1})),$
@@
where in (a) we note that component ${\boldsymbol{o}}$ of the state is initialized as a zero vector and then deterministically evolves according to $\sigma$; any sequence of ${\boldsymbol{o}}$-vectors that does not evolve according to $\sigma$ has zero probability under probability measure $\mathbb{P}_{\pi_{\text{O}}}^{\text{O}}$. In (b) we used the fact that any stationary policy $\pi_{\text{O}}\in\Pi_{\text{S}}$ for $\mathcal{M}_{\text{O}}$ can be mapped to a particular non-Markovian policy $\pi\in\Pi_{\text{NM}}$ in $\mathcal{M}_{f}$. In (c) we recall that, for $\omega=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H-1},a_{H-1})$, $\mathbb{P}_{\pi}\left[\omega\right]=p_{0}(s_{0})\cdot\pi(a_{0}|h_{0})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\pi(a_{1}|h_{1})\cdot P^{a_{1}}(s_{1},s_{2})\ldots P^{a_{H-2}}(s_{H-2},s_{H-1})\cdot\pi(a_{H-1}|h_{H-1})$.
@@
$=\sum_{\omega_{\text{O}}\in\Omega_{\text{O}}}\mathbb{P}_{\pi_{\text{O}}}^{\text{O}}[\omega_{\text{O}}]c_{\text{O}}(\{s_{H},{\boldsymbol{o}}_{H}\})$
$=\sum_{\omega_{\text{O}}\in\Omega_{\text{O}}}\mathbb{P}_{\pi}\left[\omega\right]\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\boldsymbol{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\boldsymbol{o}}_{1}=\sigma(s_{0},{\boldsymbol{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\mathbf{1}({\boldsymbol{o}}_{H}=\sigma(s_{H-1},{\boldsymbol{o}}_{H-1},a_{H-1}))f\left(\frac{1-\gamma}{1-\gamma^{H}}{\boldsymbol{o}}_{H}\right)$
=$\overset{(a)}{=}\sum_{\omega_{\text{O}}\in\Omega_{\text{O}}}\mathbb{P}_{\pi}\left[\omega\right]\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\boldsymbol{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\boldsymbol{o}}_{1}=\sigma(s_{0},{\boldsymbol{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\mathbf{1}({\boldsymbol{o}}_{H}=\sigma(s_{H-1},{\boldsymbol{o}}_{H-1},a_{H-1}))f\left(\mathbf{d}_{\omega}^{\pi,H}\right)$
=$\overset{(b)}{=}\sum_{\omega\in\Omega}\mathbb{P}_{\pi}\left[\omega\right]f\left(\mathbf{d}^{\pi,H}_{\omega}\right)\sum_{{\boldsymbol{o}}_{0},{\boldsymbol{o}}_{1},\ldots,{\boldsymbol{o}}_{H}\in\mathcal{O}}\sum_{s_{H}\in\mathcal{S}}P^{a_{H-1}}(s_{H-1},s_{H})\cdot$
$\qquad\qquad\mathbf{1}({\boldsymbol{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\boldsymbol{o}}_{1}=\sigma(s_{0},{\boldsymbol{o}}_{0},a_{0}))\cdot\ldots\cdot\mathbf{1}({\boldsymbol{o}}_{H}=\sigma(s_{H-1},{\boldsymbol{o}}_{H-1},a_{H-1}))$
=$\overset{(c)}{=}\sum_{\omega\in\Omega}\mathbb{P}_{\pi}\left[\omega\right]f\left(\mathbf{d}^{\pi,H}_{\omega}\right)$
@@
where in (a) we noted that, for any $\omega_{\text{O}}\in\Omega_{\text{O}}$, $f\left(\frac{1-\gamma}{1-\gamma^{H}}{\boldsymbol{o}}_{H}\right)=f\left(\mathbf{d}_{\omega}^{\pi,H}\right)$. In (b), we split the sum over $\omega_{\text{O}}\in\Omega_{\text{O}}$ as a sum over $\omega\in\Omega$, a sum over each possible vector ${\boldsymbol{o}}\in\mathcal{O}$ across all timesteps, and a sum over the final state $s_{H}\in\mathcal{S}$ (not included in $\omega$). We also rearranged the sums by noting that some terms do not depend on some of the sums. In (c) we note that the inner sums over the ${\boldsymbol{o}}$-vectors and $s_{H}$ equal one.
@@
$\mathcal{R}_{H}(\pi)=F_{1,H}(\pi)-\min_{\pi_{H}\in\Pi_{\textnormal{NM}}}\left\{F_{1,H}(\pi_{H})\right\}=J_{\text{O}}(\pi_{\text{O}})-\min_{\pi^{\prime}_{\text{O}}\in\Pi_{\text{S}}}J_{\text{O}}(\pi^{\prime}_{\text{O}}),$

### v2

@@ B.4 Proof of Proposition 2
There exists a one-to-one mapping between histories $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{l}$ in $\mathcal{M}_{f}$, with $0\leq l\leq H-1$, and states $\{s,{\bm{o}}\}\in\mathcal{S}\times\mathcal{O}$ in $\mathcal{M}_{\text{O}}$.
@@
For a given history $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})\in\mathcal{S}\times(\mathcal{S}\times\mathcal{A})^{l}$ in $\mathcal{M}_{f}$, with $0\leq l\leq H-1$, consider the mapping defined below that associates $h_{l}$ to a given state $\{s,{\bm{o}}\}\in\mathcal{S}\times\mathcal{O}$ for $\mathcal{M}_{\text{O}}$ by letting
@@
We aim to show that the mapping above is a bijection between the set of possible histories in $\mathcal{M}_{f}$ and the discrete state space $\mathcal{O}$ in $\mathcal{M}_{\text{O}}$. Clearly, from the mapping above defined, each history $h_{l}$ in $\mathcal{M}_{f}$ is associated with a unique state in $\mathcal{M}_{\text{O}}$. Thus, what remains is to show that any two states $\{s_{1},{\bm{o}}_{1}\}$ and $\{s_{2},{\bm{o}}_{2}\}$ for $\mathcal{M}_{\text{O}}$ are equal under mapping equation 15 if and only if their associated histories $h^{1}$ and $h^{2}$ are equal. We now make two observations. First, for a given state $\{s,{\bm{o}}\}$ in $\mathcal{M}_{\text{O}}$, component $s$ is directly related, through mapping equation 15, to the last state in the history $h_{l}$. Second, each history $h_{l}=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{l})$ will yield through mapping equation 15 a running occupancy $\bm{o}$} satisfying $\sum_{s,a}o(s,a)=\frac{1-\gamma^{l}}{1-\gamma}$; thus, histories $h_{l}$ with different lengths will yield different $\bm{o}$}-vectors. Hence, we only need to show that two running occupancies ${\bm{o}}_{1}$ and ${\bm{o}}_{2}$, associated with histories $h^{1}$ and $h^{2}$ (both of length $l$), respectively, are the same if and only if their histories up to timestep l1l-1 are the same:
  • If two histories $h^{1}$ and $h^{2}$ of length $l$ are the same, then it should be clear that their respective running occupancies, as defined through equation 15, are also the same.
  • If two running occupancies ${\bm{o}}_{1}$ and ${\bm{o}}_{2}$ are the same, then their associated histories are also the same. To prove this implication, we focus our attention to a given entry $(s,a)$ of the vectors ${\bm{o}}_{1}$ and ${\bm{o}}_{2}$. Running occupancy ${\bm{o}}_{1}$ is associated with an arbitrary history $h^{1}=(s_{0}^{1},a_{0}^{1},s_{1}^{1},a_{1}^{1},\ldots,s_{l}^{1})$; running occupancy ${\bm{o}}_{2}$ is associated with an arbitrary history $h^{2}=(s_{0}^{2},a_{0}^{2},s_{1}^{2},a_{1}^{2},\ldots,s_{l}^{2})$. If ${\bm{o}}_{1}={\bm{o}}_{2}$ then, for any $s\in\mathcal{S},a\in\mathcal{A}$,
@@
Thus, we conclude that there exists a one-to-one mapping, as defined in equation 15, between every possible history in $\mathcal{M}_{f}$ up to timestep H1H-1 and states in $\mathcal{M}_{\text{O}}$. ∎
@@ B.5 Proof of Theorem 2
The problem of finding a policy $\pi\in\Pi_{\textnormal{NM}}$ satisfying $\mathrm{OptGap}_{H}(\pi)\leq\epsilon$, for any $\epsilon\in\mathbb{R}_{0}^{+}$, can be reduced to the problem of finding a policy $\pi_{\text{O}}\in\Pi_{\textnormal{S}}$ satisfying $J_{\text{O}}(\pi_{\text{O}})-J_{\text{O}}^{*}\leq\epsilon$. In particular, if $\pi^{*}_{\text{O}}=\operatorname*{arg\,min}_{\pi_{\text{O}}\in\Pi_{\textnormal{S}}}J_{\text{O}}(\pi_{\text{O}})$, then the corresponding non-Markovian policy $\pi$ in $\mathcal{M}_{f}$ satisfies $\mathrm{OptGap}_{H}(\pi)=0$. Finally, it holds that $\mathrm{OptGap}_{H}(\pi)=J_{\text{O}}(\pi_{\text{O}})-J^{*}_{\text{O}}$, where $\pi_{O}$ is the stationary policy for $\mathcal{M}_{\text{O}}$ associated with the non-Markovian policy $\pi$ for $\mathcal{M}_{f}$.
@@
for $F_{1,H}(\pi)$ as defined in equation 6 and $J_{\text{O}}(\pi_{\text{O}})$ as defined in equation 9, where $\pi_{\text{O}}$ is the stationary policy for $\mathcal{M}_{\text{O}}$ associated with the non-Markovian policy $\pi$ for $\mathcal{M}_{f}$.
@@
For any $H\in\mathbb{N}$, finite-horizon random trajectories $(\mathrm{s}_{0},\mathrm{a}_{0},\mathrm{s}_{1},\mathrm{a}_{1},\ldots,\mathrm{s}_{H-1},\mathrm{a}_{H-1})$ in $\mathcal{M}_{f}$ are associated with the probability space $(\Omega,\mathcal{F},\mathbb{P}_{\pi})$. We write specific trajectories as $\omega\in\Omega$, with $\omega=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H-1},a_{H-1})$. We highlight that the probability of a given trajectory $\omega\in\Omega$ under policy $\pi\in\Pi_{\text{NM}}$ can be calculated as $\mathbb{P}_{\pi}\left[\omega\right]=p_{0}(s_{0})\cdot\pi(a_{0}|h_{0})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\pi(a_{1}|h_{1})\cdot P^{a_{1}}(s_{1},s_{2})\ldots P^{a_{H-2}}(s_{H-2},s_{H-1})\cdot\pi(a_{H-1}|h_{H-1})$. On the other hand, random trajectories $(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\mathrm{a}_{0},\{\mathrm{s}_{1},\mathbf{o}_{1}\},\mathrm{a}_{1},\ldots,\{\mathrm{s}_{H},\mathbf{o}_{H}\})$ in $\mathcal{M}_{\text{O}}$ are associated with probability space $(\Omega_{\text{O}},\mathcal{F}_{\text{O}},\mathbb{P}_{\pi_{\text{O}}}^{O})$. We write specific trajectories as $\omega_{\text{O}}\in\Omega_{\text{O}}$, with $\omega_{\text{O}}=(\{s_{0},{\bm{o}}_{0}\},a_{0},\{s_{1},{\bm{o}}_{1}\},a_{1},\ldots,\{s_{H},{\bm{o}}_{H}\})$.
We start by noting that, for any trajectory $\omega_{\text{O}}=(\{s_{0},{\bm{o}}_{0}\},a_{0},\{s_{1},{\bm{o}}_{1}\},a_{1},\ldots,\{s_{H},{\bm{o}}_{H}\})\in\Omega_{\text{O}}$,
$\mathbb{P}_{\pi_{\text{O}}}^{\text{O}}[\omega_{\text{O}}]$ $=p_{0,\text{O}}(\{s_{0},{\bm{o}}_{0}\})\cdot\pi_{\text{O}}(a_{0}|\{s_{0},{\bm{o}}_{0}\})\cdot P_{\text{O}}^{a_{0}}(\{\mathrm{s}_{0},\mathbf{o}_{0}\},\{\mathrm{s}_{1},\mathbf{o}_{1}\})\cdot\ldots$
$\qquad\qquad\cdot\pi_{\text{O}}(a_{H-1}|\{s_{H-1},{\bm{o}}_{H-1}\})\cdot P_{\text{O}}^{a_{H-1}}(\{s_{H-1},{\bm{o}}_{H-1}\},\{s_{H},{\bm{o}}_{H}\}).$
$\overset{(a)}{=}p_{0}(s_{0})\cdot\mathbf{1}({\bm{o}}_{0}=[0,\ldots,0])\cdot\pi_{\text{O}}(a_{0}|\{s_{0},{\bm{o}}_{0}\})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\mathbf{1}({\bm{o}}_{1}=\sigma(s_{0},{\bm{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\pi_{\text{O}}(a_{H-1}|\{s_{H-1},{\bm{o}}_{H-1}\})\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\bm{o}}_{H}=\sigma(s_{H-1},{\bm{o}}_{H-1},a_{H-1}))$
$\overset{(b)}{=}p_{0}(s_{0})\cdot\mathbf{1}({\bm{o}}_{0}=[0,\ldots,0])\cdot\pi(a_{0}|h_{0})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\mathbf{1}({\bm{o}}_{1}=\sigma(s_{0},{\bm{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\pi(a_{H-1}|h_{H-1})\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\bm{o}}_{H}=\sigma(s_{H-1},{\bm{o}}_{H-1},a_{H-1}))$
@@
$\overset{(c)}{=}\mathbb{P}_{\pi}\left[\omega\right]\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\bm{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\bm{o}}_{1}=\sigma(s_{0},{\bm{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\mathbf{1}({\bm{o}}_{H}=\sigma(s_{H-1},{\bm{o}}_{H-1},a_{H-1})),$
@@
where in (a) we note that component $\bm{o}$} of the state is initialized as a zero vector and then deterministically evolves according to $\sigma$; any sequence of $\bm{o}$}-vectors that does not evolve according to $\sigma$ has zero probability under probability measure $\mathbb{P}_{\pi_{\text{O}}}^{\text{O}}$. In (b) we used the fact that any stationary policy $\pi_{\text{O}}\in\Pi_{\text{S}}$ for $\mathcal{M}_{\text{O}}$ can be mapped to a particular non-Markovian policy $\pi\in\Pi_{\text{NM}}$ in $\mathcal{M}_{f}$. In (c) we recall that, for $\omega=(s_{0},a_{0},s_{1},a_{1},\ldots,s_{H-1},a_{H-1})$, $\mathbb{P}_{\pi}\left[\omega\right]=p_{0}(s_{0})\cdot\pi(a_{0}|h_{0})\cdot P^{a_{0}}(s_{0},s_{1})\cdot\pi(a_{1}|h_{1})\cdot P^{a_{1}}(s_{1},s_{2})\ldots P^{a_{H-2}}(s_{H-2},s_{H-1})\cdot\pi(a_{H-1}|h_{H-1})$.
@@
$=\sum_{\omega_{\text{O}}\in\Omega_{\text{O}}}\mathbb{P}_{\pi_{\text{O}}}^{\text{O}}[\omega_{\text{O}}]c_{\text{O}}(\{s_{H},{\bm{o}}_{H}\})$
$=\sum_{\omega_{\text{O}}\in\Omega_{\text{O}}}\mathbb{P}_{\pi}\left[\omega\right]\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\bm{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\bm{o}}_{1}=\sigma(s_{0},{\bm{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\mathbf{1}({\bm{o}}_{H}=\sigma(s_{H-1},{\bm{o}}_{H-1},a_{H-1}))f\left(\frac{1-\gamma}{1-\gamma^{H}}{\bm{o}}_{H}\right)$
$\overset{(a)}{=}\sum_{\omega_{\text{O}}\in\Omega_{\text{O}}}\mathbb{P}_{\pi}\left[\omega\right]\cdot P^{a_{H-1}}(s_{H-1},s_{H})\cdot\mathbf{1}({\bm{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\bm{o}}_{1}=\sigma(s_{0},{\bm{o}}_{0},a_{0}))\cdot\ldots$
$\qquad\qquad\cdot\mathbf{1}({\bm{o}}_{H}=\sigma(s_{H-1},{\bm{o}}_{H-1},a_{H-1}))f\left(\mathbf{d}_{\omega}^{\pi,H}\right)$
$\overset{(b)}{=}\sum_{\omega\in\Omega}\mathbb{P}_{\pi}\left[\omega\right]f\left(\mathbf{d}^{\pi,H}_{\omega}\right)\sum_{{\bm{o}}_{0},{\bm{o}}_{1},\ldots,{\bm{o}}_{H}\in\mathcal{O}}\sum_{s_{H}\in\mathcal{S}}P^{a_{H-1}}(s_{H-1},s_{H})\cdot$
$\qquad\qquad\mathbf{1}({\bm{o}}_{0}=[0,\ldots,0])\cdot\mathbf{1}({\bm{o}}_{1}=\sigma(s_{0},{\bm{o}}_{0},a_{0}))\cdot\ldots\cdot\mathbf{1}({\bm{o}}_{H}=\sigma(s_{H-1},{\bm{o}}_{H-1},a_{H-1}))$
$\overset{(c)}{=}\sum_{\omega\in\Omega}\mathbb{P}_{\pi}\left[\omega\right]f\left(\mathbf{d}^{\pi,H}_{\omega}\right)$
@@
where in (a) we noted that, for any $\omega_{\text{O}}\in\Omega_{\text{O}}$, $f\left(\frac{1-\gamma}{1-\gamma^{H}}{\bm{o}}_{H}\right)=f\left(\mathbf{d}_{\omega}^{\pi,H}\right)$. In (b), we split the sum over $\omega_{\text{O}}\in\Omega_{\text{O}}$ as a sum over $\omega\in\Omega$, a sum over each possible vector ${\bm{o}}\in\mathcal{O}$ across all timesteps, and a sum over the final state $s_{H}\in\mathcal{S}$ (not included in $\omega$). We also rearranged the sums by noting that some terms do not depend on some of the sums. In (c) we note that the inner sums over the $\bm{o}$}-vectors and $s_{H}$ equal one.
@@
$\mathrm{OptGap}_{H}(\pi)=F_{1,H}(\pi)-\min_{\pi_{H}\in\Pi_{\textnormal{NM}}}\left\{F_{1,H}(\pi_{H})\right\}=J_{\text{O}}(\pi_{\text{O}})-\min_{\pi^{\prime}_{\text{O}}\in\Pi_{\text{S}}}J_{\text{O}}(\pi^{\prime}_{\text{O}}),$